OpenStudy (anonymous):

3. Marvel Man movie is being released to a limited number of movie theaters. The number of showtimes for the movie is twice the number of movie theaters. The number of seats available for each showtime is three times the number of showtimes. Let x represent the number of theatres showing the movie. 1. Determine the three factors that are to be multiplied to determine the total number of seats available for the movie. 2. When simplified what is the degree and classification of this expression? 3. Interpet each pairing of factors as a single unit in the context of the problem

OpenStudy (ranga):

Number of theaters showing the movie = x "number of showtimes for the movie is twice the number of movie theaters" Number of showtimes for the movie = 2 * x "number of seats available for each showtime is three times the number of showtimes" 1) Number of seats available for each showtime = 3 * 2 * x = 6x The three factors that need to be multiplied to determine the TOTAL number of seats are: 6x, 2x and x.